One number is two more than five times another. their product is 24. Find the numbers.
Answer by farah(18)

xy=24,5x+2=y,,so x(5x+2)=24,5x^2+2x=24,5x^2+2x-24=0,find the root n u'll get x=2 n -24,take the positive num which is 2,so if xy=24,when x=2,y=12,hence the number is 2 and 12,sorry if i'm doing wrong.